# Articulatory Encodec: Coding Speech through Vocal Tract Kinematics

### v2

@@ Abstract
Vocal tract articulation is a natural, grounded control space of speech production. The spatiotemporal coordination of articulators combined with the vocal source shapes intelligible speech sounds to enable effective spoken communication. Based on this physiological grounding of speech, we propose a new framework of neural encoding-decoding of speech – Articulatory Encodec. Articulatory Encodec comprises an articulatory analysis model that infers articulatory features from speech audio, and an articulatory synthesis model that synthesizes speech audio from articulatory features. The articulatory features are kinematic traces of vocal tract articulators and source features, which are intuitively interpretable and controllable, being the actual physical interface of speech production. An additional speaker identity encoder is jointly trained with the articulatory synthesizer to inform the voice texture of individual speakers. By training on large-scale speech data, we achieve a fully intelligible, high-quality articulatory synthesizer that generalizes to unseen speakers. Furthermore, the speaker embedding is effectively disentangled from articulations, which enables accent-perserving zero-shot voice conversion. To the best of our knowledge, this is the first demonstration of universal, high-performance articulatory inference and synthesis, suggesting the proposed framework as a powerful coding system of speech.
@@
neural codec, speech coding, speech synthesis, articulatory synthesis, speech inversion, acoustics-to-articulatory inversion, electromagnetic articulography.
@@ I Introduction
Humans naturally produce intelligible speech by controlling articulators on the vocal tract. Such vocal tract articulation has long been claimed to be the physiological ground of speech production in various aspects. The source-filter theory of speech describes articulation as shaping the vocal cavity to implement filters on source, or glottal flow, to create speech sounds [1, 2]. Articulatory phonetics and phonology have explained the basis of speech in terms of the coordination of articulators, identifying some canonical articulators that can determine the phonetic properties [3, 4, 5]. In cognitive neuroscience, the speech sensorimotor cortex has been proven to represent continuous, real-time vocal tract articulation while naturally speaking, suggesting the vocal tract articulation as a cognitive basis of speech production [6, 7, 8].
@@
Furthermore, the recent findings by Cho et al. [9, 10] suggest that articulatory inversion naturally emerges from self-supervised learning (SSL) of speech. When probed on articulatory kinematics measured by electromagnetic articulography (EMA), the feature representation of the recent speech SSL models (e.g., HuBERT [11]) is highly correlated with EMA, where high-fidelity articulation can be reconstructed by a simple linear mapping from speech SSL features [9]. This suggests that the articulatory inference is a natural solution of SSL of speech for abstracting speech information. This emergent property is further shown to be universal to any speakers, dialects, and even languages [10]. Together, these suggest that the biophysical, articulatory representation of speech is a shared coding principle in both biological and artificial intelligence of speech. This convergence raises an interesting question – can we express any arbitrary speech using articulatory representation, or articulatory code?
Figure 1: Articulatory encodec framework.
@@
Previous studies have demonstrated that intelligible speech can be synthesized from articulatory features [12, 13, 14, 15, 16], and combined with acoustics-to-articulatory inversion (AAI), resynthesis frameworks have shown the potential of articulatory features as viable intermediate for speech coding systems [17, 16]. However, the previous methods are limited to a fixed set of speakers and the qualities are still far behind the commercial speech synthesis models. This absence of a universal, generalizable framework has significantly limited the utility of articulatory-based speech coding as a practically usable system.
@@
Here, we first demonstrate a high-performance, universal articulatory encoder and decoder that can scale and generalize across an indefinite number of speakers. We leverage the universal articulatory inference by speech SSL [10] to build a generalizable articulatory encoder that transforms speech into kinematic traces in a template articulatory space. The template articulatory space is agnostic to individual anatomical differences which are compensated by a separate speaker identity encoder. By training a synthesis model, or a vocoder, on a large-scale dataset, we achieve a universal articulatory vocoder that can generate fully intelligible, high-quality speech from any speaker’s articulation. Furthermore, the speaker identity encoder successfully disentangle speaker identity from articulation, which is demonstrated by a zero-shot, accent-preserving voice conversion. By closing the loop of articulatory encoding and decoding, we propose a novel, speech science guided encoding-decoding (encodec) framework of speech – Articulatory Encodec (Fig. 1). [^1] We will open-source the code and model checkpoints upon publication. The Articulatory Encodec framework shows a minimal loss of intelligibility and quality compared to the original speech audio. [^2] Audio samples are available at https://articulatoryencodec.github.io.
@@
  • Low-dimensionality: The articulatory features have only 14 channels with 50 Hz sampling rate. This is significantly lower than the previous acoustic features or neural embedding of speech.
@@
  • Controllability: The articulatory features can be naturally controlled by the same principle as speech production.
@@
With these unique benefits, empirical evidence and demonstration show the promising potential of Articulatory Encodec as a valid, novel coding framework of speech.
@@ II-B Acoustics-to-Articulation Inversion
To replace the complicated data collection procedure, acoustic-to-articulatory inversion (AAI) models have been actively developed to predict EMA directly from speech audio [24, 25, 26, 6, 7, 17, 27, 28, 16]. However, the individual variance in vocal tract anatomy across speakers induces inconsistent placements of sensors, which has posed a significant barrier to developing a model that can generalize to unseen speakers [29, 17, 27, 28]. Despite such variability, a canonical basis of articulation is suggested to exist, which is agnostic to individual vocal tract anatomy [4, 30, 17, 10]. In fact, Cho et al. [10] demonstrated that a linear affine transformation can geometrically align one speaker’s articulatory system to another’s. This suggests that individual articulatory spaces are lying on the same linear space so that an articulatory space of any speaker can be a hypothetical universal template space of articulation. We empirically prove this statement by leveraging a single-speaker AAI model as a universal articulatory encoder for our codec framework.
@@ II-D Neural Coding of Speech
Here, we propose articulatory features as interpretable, controllable, and grounded coding of speech that can fully represent any arbitrary speech. A similar concurrent work is proposed towards interpretable speech representations, which utilizes sparse phonetic posteriorgams instead of articulatory features [53].
Figure 2: Pipeline of articulatory analysis and vocoder. The articulatory analysis is composed of vocal tract articulation, source features, and speaker embedding, which are then fed to the synthesizer (HiFi-GAN) in the synthesis pipeline. The modules colored with orange are updated while training the vocoder and other modules are fixed.
@@ III-A Articulatory Analysis
In Articulatory Encodec, speech is encoded as factors obtained by an analytic framework that infers three different components of speech: vocal tract articulation, source features, and speaker identity (Fig. 2). The first two are based on pre-trained analysis models that provide kinematic traces of the physical articulators, and on vocal source features. The last component, speaker identity, is inferred by a model jointly trained with the synthesizer.
@@ III-A1 Vocal Tract Articulation
The second hypothesis is that the SSL-linear AAI is speaker agnostic. This again requires the following propositions to hold: there exists a subspace in SSL features that is agnostic to voice identity, and a linear mapping used to project the features to the subspace is also speaker independent. This means that even if the SSL-linear AAI is trained only on the MNGU0 speaker’s voice, the model should project speech from other speakers to the same articulatory space. We empirically prove this hypothesis in §IV-D, where the inference remains consistent with different voice identities.
@@ IV Experiments & Results
We evaluate the proposed articulatory encodec in different aspects to validate each part of the framework – universality of articulatory encoding (IV-A), information conservation in encoding-decoding (IV-B), multilingual generalization (IV-C), and disentanglement of speaker identity encoding (IV-D).
@@ IV-A Universality of Single-Speaker Articulatory Encoding
IV-B Performance of Resynthesis by Articulatory Encodec
TABLE II: Performance comparison of ground truth speech (GT) and resynthesized speech by articulatory encodec (Resynth)
@@ IV-B Performance of Resynthesis by Articulatory Encodec
Furthermore, the results on VCTK also demonstrate high intelligibility with WER of 3.73% and CER of 1.96%. Though the gap from the ground truth is larger than LibriTTS-R, this is a remarkable level of performance given that the model is not exposed to the entire dataset during training. As LibriTTS-R has cleaner audio than VCTK, the model demonstrates some speech enhancement capacity which results in a higher MOS and UTMOS than the ground truth. While the speakers in LibriTTS-R are concentrated on American English, VCTK includes a variety of accents, primarily focusing on English speakers with different regional accents from the United Kingdom. Therefore, the high performance demonstrated on VCTK suggests that Articulatory Encodec can robustly generalize to unseen speakers and accents.
@@ IV-C Multilingual Generalization
We evaluate Articulatory Encodec on speech corpora from other languages: 7 European languages (German, Dutch, Portuguese, Italian, Polish, Spanish, French) from multilingual LibriSpeech [70] and 3 East Asian languages (Korean, Japanese, and Chinese (Mandarin)) from KSS [71], JVS [72], and AISHELL [73], respectively. Except for KSS, all corpora include multiple speakers. We randomly sampled 200 utterances from the test split of each corpus to evaluate the resynthesized speech. [^4] We excluded samples with numerals to avoid converting Arabic numerals. Table III compares the intelligibility (WER and CER using Whisper) and quality (UTMOS). As no interword spacing is used for Japanese and Chinese orthographic systems, words in these languages are not separated as straightforwardly as other languages. Therefore, we omit the WER for these languages. We also evaluate the performance of a multilingual version of the model that is obtained by fine-tuning to the multilingual data (denoted as ML-R in Table III). The original English-only trained encodec is denoted as EN-R. See Appendix B.8. for details of the fine-tuning procedure.
@@ IV-D Speaker Recognition and Voice Control by Speaker Identity
The high coding-recoding similarity in articulation suggests that the articulation is well disentangled from speaker identity and pitch, which indicates that accent is preserved after the voice conversion. The correlation is as high as 0.944 and is marginally different from the self-targeted resynthesis, 0.954, which also indicates the stability of the coding by our articulatory encodec. This result further supports the claim that the proposed AAI is agnostic to speaker identity, yielding highly consistent inferences even with different voice identities. Moreover, the similarities outperform baselines, especially in pitch, suggesting our model better disentangles pitch from the other features. Lastly, the intelligibility slightly decreases after the voice conversion, but the difference is marginal and the speech remains highly intelligible with WER of 4.83 %. Both baselines shows lower WERs than our model (FreeVC: 4.00 %, and QuickVC: 4.17 %0) indicating a potential internal mechanism that might improve intelligibility by modifying the articulatory features of the original speech. Finally, the voice conversion shows some degraded UTMOS as 3.83, 0.17 decrease compared to the resynthesis, and QuickVC shows the highest UTMOS as 4.22. This suggests that our model has a potential trade-off between disentanglement and audio quality of the synthesized audio. However, this degrades in naturalness may be confounded by a natural correlation between accents and voice textures, so that some voice conversion cases may result in rare combinations of accent and speaker identity, harming the UTMOS score. To sum up, these results indicate that Articulatory Encodec is a stable and disentangled coding system of speech.
@@ V-A Case 1: Place of Articulation – “lock” vs “rock”
We demonstrate an articulatory control simulation by interpolating tongue articulations between the words [14]. The mixing factor, $\alpha$, is applied to weigh the sum of each of the tongue articulatory traces (TT, TB, and TD), i.e., $Art("lock")\cdot\alpha+Art("rock")\cdot(1-\alpha)$, where $Art(\cdot)$ means articulatory encoding. The Fig. 5 bottom panel shows the interpolated traces with $\alpha\in\{1.0,0.8,0.6,0.4,0.2,0.0,-0.2\}$, gradually changing pronunciation from ”lock” to ”rock”. Then, ASR is applied to synthesized speech to determine which word is perceived, drawing a perceptual boundary at $\alpha=0.2$, which is denoted using different line styles in Fig. 5. This demonstration shows that Articulatory Encodec provides interpretable control knobs, which can be manipulated to simulate speech sounds and observe the causal interaction between articulatory control and the generated speech.
@@ V-B Case 2: Manner of Articulation – Voice Onset Time
These two proof-of-concept showcases suggest the promising utility of Articulatory Encodec as a physical simulator of articulatory control, and as a speech analysis tool that provides a natural and intuitive phonological interpretation.
@@ V-C Potential Applications
Based on the results, we envision several potential applications of Articulatory Encodec. First, the framework can be used as an analysis platform for investigating a phonological basis of speech without collecting high-cost articulatory data [16]. Second, the high-performance synthesizer can be used as a speech simulator, which can facilitate a control theoretic approach to speech processing or reinforcement learning of speaking agents [3, 81, 41, 82]. Third, the universal articulatory encoding can be utilized for a language learning tool or therapy, where visualizing the vocal tract can enhance the learning experience [83, 84, 85]. Lastly, a TTS system can be built upon Articulatory Encodec that synthesizes articulatory control from text or higher-order structure of speech [86, 87, 88], revealing a descriptive and interpretable relationship between text and articulation. In conclusion, our framework holds promise for various applications in both speech science and engineering.

### v3

# Coding Speech through Vocal Tract Kinematics

@@ Abstract
Vocal tract articulation is a natural, grounded control space of speech production. The spatiotemporal coordination of articulators combined with the vocal source shapes intelligible speech sounds to enable effective spoken communication. Based on this physiological grounding of speech, we propose a new framework of neural encoding-decoding of speech – Speech Articulatory Coding (SPARC). SPARC comprises an articulatory analysis model that infers articulatory features from speech audio, and an articulatory synthesis model that synthesizes speech audio from articulatory features. The articulatory features are kinematic traces of vocal tract articulators and source features, which are intuitively interpretable and controllable, being the actual physical interface of speech production. An additional speaker identity encoder is jointly trained with the articulatory synthesizer to inform the voice texture of individual speakers. By training on large-scale speech data, we achieve a fully intelligible, high-quality articulatory synthesizer that generalizes to unseen speakers. Furthermore, the speaker embedding is effectively disentangled from articulations, which enables accent-perserving zero-shot voice conversion. To the best of our knowledge, this is the first demonstration of universal, high-performance articulatory inference and synthesis, suggesting the proposed framework as a powerful coding system of speech.
@@
speech coding, speech synthesis, articulatory synthesis, speech inversion, acoustics-to-articulatory inversion, electromagnetic articulography.
@@ I Introduction
Humans naturally produce intelligible speech by controlling articulators on the vocal tract. Such vocal tract articulation has long been claimed to be the physiological ground of speech production in various aspects [1]. The source-filter theory of speech describes articulation as shaping the vocal cavity to implement filters on source, or glottal flow, to create speech sounds [2, 3]. Articulatory phonetics and phonology have explained the basis of speech in terms of the coordination of articulators, identifying some canonical articulators that can determine the phonetic properties [4, 1, 5]. In cognitive neuroscience, the speech sensorimotor cortex has been proven to represent continuous, real-time vocal tract articulation while naturally speaking, suggesting the vocal tract articulation as a cognitive basis of speech production [6, 7, 8].
@@
Furthermore, the recent findings by Cho et al. [9, 10] suggest that articulatory inversion naturally emerges from self-supervised learning (SSL) of speech. When probed on articulatory kinematics measured by electromagnetic articulography (EMA), the feature representation of the recent speech SSL models (e.g., HuBERT [11]) is highly correlated with EMA, where high-fidelity articulation can be reconstructed by a simple linear mapping from speech SSL features [9]. This suggests that the articulatory inference is a natural solution of SSL of speech for abstracting speech information. This emergent property is further shown to be universal across speakers, dialects, and even languages [10]. Together, these suggest that the biophysical, articulatory representation of speech is a shared coding principle in both biological and artificial intelligence of speech. This convergence raises an interesting question – Can we represent any arbitrary speech using articulatory features?
Figure 1: The Speech Articulatory Coding (SPARC) framework. It encodes speech using articulatory features (analysis) and decodes these features back into speech (synthesis).
@@
Previous studies have demonstrated that intelligible speech can be synthesized from articulatory features [12, 13, 14, 15, 16], and combined with acoustics-to-articulatory inversion (AAI), resynthesis frameworks have shown the potential of articulatory features as viable intermediate for speech coding systems [17, 16]. However, these previous methods are limited to a fixed set of speakers and the quality is still far behind the commercial speech synthesis models. This absence of a universal, generalizable framework has significantly limited the practical utility of articulatory-based speech coding.
@@
Here, we first demonstrate a high-performance, universal articulatory encoder and decoder that can scale and generalize across an indefinite number of speakers. We leverage the universal articulatory inference by speech SSL [10] to build a generalizable articulatory encoder that transforms speech into vocal tract kinematic traces in a template articulatory space. The template articulatory space is agnostic to individual anatomical differences which are compensated by a separate speaker identity encoder. By training a synthesis model, or a vocoder, on a large-scale dataset, we achieve a universal articulatory vocoder that can generate fully intelligible, high-quality speech from any speaker’s articulation. Furthermore, the speaker identity encoder successfully disentangles speaker identity from articulation, which is demonstrated by a zero-shot, accent-preserving voice conversion. By closing the loop of articulatory encoding and decoding, we propose a novel, speech science guided encoding-decoding framework of speech – Speech Articulatory Coding (SPARC) (Fig. 1). [^1] The code is available at https://github.com/Berkeley-Speech-Group/Speech-Articulatory-Coding. The SPARC framework shows a minimal loss of intelligibility and quality compared to the original speech audio. [^2] Audio samples: https://berkeley-speech-group.github.io/sparc-demo
@@
  • Low-dimensionality: The articulatory features have only 14 channels with 50 Hz sampling rate.
@@
  • Controllability: The features can be naturally controlled by the same principle as speech production.
@@
With these unique benefits, we demonstrates empirical evidence of the promising potential of SPARC as a valid, novel coding framework for speech.
@@ II-B Acoustics-to-Articulation Inversion
To replace the complicated data collection procedure, acoustic-to-articulatory inversion (AAI) models have been actively developed to predict EMA directly from speech audio [24, 25, 26, 6, 7, 17, 27, 28, 16]. However, the individual variance in vocal tract anatomy across speakers induces inconsistent placements of sensors, which has posed a significant barrier to developing a model that can generalize to unseen speakers [29, 17, 27, 28]. Despite such variability, a canonical basis of articulation is suggested to exist, which is agnostic to individual vocal tract anatomy [1, 30, 17, 10]. In fact, Cho et al. [10] demonstrated that a linear affine transformation can geometrically align one speaker’s articulatory system to another’s. This suggests that individual articulatory spaces are lying on the same linear space so that an articulatory space of any speaker can be a hypothetical universal template space of articulation. We empirically validate this statement by using a single-speaker AAI model as a universal articulatory encoder for our coding framework.
@@ II-D Neural Coding of Speech
Here, we propose articulatory features as interpretable, controllable, and grounded coding of speech that can fully represent any arbitrary speech. A similar concurrent work is proposed towards interpretable speech representations, which utilizes sparse phonetic posteriorgrams instead of articulatory features [53].
Figure 2: Pipeline of articulatory analysis and synthesis. The articulatory analysis is composed of vocal tract articulation, source features, and speaker embedding, which are then fed to the synthesizer (HiFi-GAN) in the synthesis pipeline. The modules colored with orange (FFN and HiFi-GAN) are updated while training the synthesis model and other modules are fixed.
@@ III-A Articulatory Analysis
In SPARC, speech is encoded as factors obtained by an analytic framework that infers three different components of speech: vocal tract articulation, source features, and speaker identity (Fig. 2). The first two are based on pre-trained analysis models that provide kinematic traces of the physical articulators, and on vocal source features. The last component, speaker identity, is inferred by a model jointly trained with the synthesizer.
@@ III-A1 Vocal Tract Articulation
The second hypothesis is that the SSL-linear AAI is speaker agnostic. This again requires the following propositions to hold: there exists a subspace in SSL features that is agnostic to voice identity, and a linear mapping used to project the features to the subspace is also speaker independent. This means that even if the SSL-linear AAI is trained only on the MNGU0 speaker’s voice, the model should project speech from other speakers to the same articulatory space. We provide empirical evidence to support this hypothesis in a later section (§IV-D), where we show that the inference remains consistent after converting between different voice identities.
@@ IV Experiments & Results
We evaluate the proposed SPARC in different aspects to validate each part of the framework – universality of articulatory encoding (IV-A), information conservation in encoding-decoding (IV-B), multilingual generalization (IV-C), and disentanglement of speaker identity encoding (IV-D).
@@ IV-A Universality of Single-Speaker Articulatory Encoding
IV-B Performance of Resynthesis by SPARC
TABLE II: Performance comparison of ground truth speech (GT) and resynthesized speech by SPARC (Resynth)
@@ IV-B Performance of Resynthesis by SPARC
Furthermore, the results on VCTK also demonstrate high intelligibility with WER of 3.73% and CER of 1.96%. Though the gap from the ground truth is larger than LibriTTS-R, this is a remarkable level of performance given that the model is not exposed to the entire dataset during training. As LibriTTS-R has cleaner audio than VCTK, the model demonstrates some speech enhancement capacity which results in a higher MOS and UTMOS than the ground truth. While the speakers in LibriTTS-R are concentrated on American English, VCTK includes a variety of accents, primarily focusing on English speakers with different regional accents from the United Kingdom. Therefore, the high performance demonstrated on VCTK suggests that SPARC can robustly generalize to unseen speakers and accents.
@@ IV-C Multilingual Generalization
We evaluate SPARC on speech corpora from other languages: 7 European languages (German, Dutch, Portuguese, Italian, Polish, Spanish, French) from multilingual LibriSpeech [70] and 3 East Asian languages (Korean, Japanese, and Chinese (Mandarin)) from KSS [71], JVS [72], and AISHELL [73], respectively. Except for KSS, all corpora include multiple speakers. We randomly sampled 200 utterances from the test split of each corpus to evaluate the resynthesized speech. [^4] We excluded samples with numerals to avoid converting Arabic numerals. Table III compares the intelligibility (WER and CER using Whisper) and quality (UTMOS). As no interword spacing is used for Japanese and Chinese orthographic systems, words in these languages are not separated as straightforwardly as other languages. Therefore, we omit the WER for these languages. We also evaluate the performance of a multilingual version of the model that is obtained by fine-tuning to the multilingual data (denoted as ML-R in Table III). The original English-only trained model is denoted as EN-R. See Appendix B.8. for details of the fine-tuning procedure.
@@ IV-D Speaker Recognition and Voice Control by Speaker Identity
The high coding-recoding similarity in articulation suggests that the articulation is well disentangled from speaker identity and pitch, which indicates that accent is preserved after the voice conversion. The correlation is as high as 0.944 and is marginally different from the self-targeted resynthesis, 0.954, which also indicates the stability of the coding by our SPARC. This result further supports the claim that the proposed AAI is agnostic to speaker identity, yielding highly consistent inferences even with different voice identities. Moreover, the similarities outperform baselines, especially in pitch, suggesting our model better disentangles pitch from the other features. Lastly, the intelligibility slightly decreases after the voice conversion, but the difference is marginal and the speech remains highly intelligible with WER of 4.83 %. Both baselines shows lower WERs than our model (FreeVC: 4.00 %, and QuickVC: 4.17 %0) indicating a potential internal mechanism that might improve intelligibility by modifying the articulatory features of the original speech. Finally, the voice conversion shows some degraded UTMOS as 3.83, 0.17 decrease compared to the resynthesis, and QuickVC shows the highest UTMOS as 4.22. This suggests that our model has a potential trade-off between disentanglement and audio quality of the synthesized audio. However, this degrades in naturalness may be confounded by a natural correlation between accents and voice textures, so that some voice conversion cases may result in rare combinations of accent and speaker identity, harming the UTMOS score. To sum up, these results indicate that SPARC is a stable and disentangled coding system of speech.
V Demonstration of Interpretability and Controllability of SPARC
@@ V-A Case 1: Place of Articulation – “lock” vs “rock”
We demonstrate an articulatory control simulation by interpolating tongue articulations between the words [14]. The mixing factor, $\alpha$, is applied to weigh the sum of each of the tongue articulatory traces (TT, TB, and TD), i.e., $Art("lock")\cdot\alpha+Art("rock")\cdot(1-\alpha)$, where $Art(\cdot)$ means articulatory encoding. The Fig. 5 bottom panel shows the interpolated traces with $\alpha\in\{1.0,0.8,0.6,0.4,0.2,0.0,-0.2\}$, gradually changing pronunciation from ”lock” to ”rock”. Then, ASR is applied to synthesized speech to determine which word is perceived, drawing a perceptual boundary at $\alpha=0.2$, which is denoted using different line styles in Fig. 5. This demonstration shows that SPARC provides interpretable control knobs, which can be manipulated to simulate speech sounds and observe the causal interaction between articulatory control and the generated speech.
@@ V-B Case 2: Manner of Articulation – Voice Onset Time
These two proof-of-concept showcases suggest the promising utility of SPARC as a physical simulator of articulatory control, and as a speech analysis tool that provides a natural and intuitive phonological interpretation.
@@ V-C Potential Applications
Based on the results, we envision several potential applications of SPARC. First, the framework can be used as an analysis platform for investigating a phonological basis of speech without collecting high-cost articulatory data [16]. Second, the high-performance synthesizer can be used as a speech simulator, which can facilitate a control theoretic approach to speech processing or reinforcement learning of speaking agents [4, 81, 41, 82]. Third, the universal articulatory encoding can be utilized for a language learning tool or therapy, where visualizing the vocal tract can enhance the learning experience [83, 84, 85]. Lastly, a TTS system can be built upon SPARC that synthesizes articulatory control from text or higher-order structure of speech [86, 87, 88], revealing a descriptive and interpretable relationship between text and articulation. In conclusion, our framework holds promise for various applications in both speech science and engineering.